OpenStudy (anonymous):

f(x)=(3x)/(x+3) and g(x)=2/(4x-1) find f of g and g of f

OpenStudy (anonymous):

\[f(g(x))\] work from the inside first \[f(\frac{2}{4x-1})\]

OpenStudy (anonymous):

then \[f(\spadesuit) = \frac{3\spadesuit}{\spadesuit +3}\] so \[f(\frac{2}{4x-1})=\frac{3\frac{2}{4x-1}}{\frac{2}{4x-1} +3}\]

OpenStudy (anonymous):

now don't I have to multiply top and bottom by (4x-1)?